3.If an edge of a cube is increased by 33%, by how much is the total surface area increased?
1
Expert's answer
2013-01-29T11:42:36-0500
Surface area of a cube S:
S = 6 a^2
a is the length of the side of each edge of the cube
If an edge of a cube is increased by 33%
a' = a + 0.33a = 1.33a
a' =& the new edge of a cube
S' = 6 a'^2 = 6 a^2 *1.33^2 = 1.33^2 *S =& 1.7689*S
S' - Surface area of a new cube
the total surface area increased in
(S'-S)/S *100% = (1.7689 - 1)*100% = 76.89%
Answer:& 76.89%